Evaluate the given expression and express the result using the usual format for writing numbers​ (instead of scientific​ notation). 52 Upper P 2
Answer #1

We asked to find 52P2 using permutation rule nP​​​​​r​​​​ = n!/(n-r)!

  

52P2 = 52!/(52-2)! = 52*51*50!/50! = 2652

So   52P2 = 2652